brandonHollinger:

Solve: negative 1 over 2 x + 1 = −x + 8 how would i do this with the fraction?

mxddi3:

\[-\frac{ 1 }{ 2 }x+1=-x+8\] so you have this right?

brandonHollinger:

but how do i find x?

mxddi3:

we're gonna get there

brandonHollinger:

ok

mxddi3:

so you see how there is a - one half x?

brandonHollinger:

yeah

mxddi3:

well, if you want to get rid of the fraction, multiply the ENTIRE thing by the denominator.this gives us \[-x+2=-2x+16\] do you get it so far?

brandonHollinger:

yes

mxddi3:

ok, so then we combine like terms. we'll start with x. since there is a -x on the left, we are going to ADD x to both sides. so what is -2x+x?

brandonHollinger:

-3x

mxddi3:

no, we have NEGATIVE two plus POSITIVE one, which equals what?

brandonHollinger:

-1x

mxddi3:

good. so we have 2=-1x+16 now, since we combines the x's, we need to do the constants. since we started on the left last time, we're going to go to the right. since we have a positive 16, we are going to subtract 16 from both sides. we now have: 2-16= -1x + 16-16 what does that give us?

brandonHollinger:

18= -1x

mxddi3:

close. you have a positive 2 and you subtract 16 from that. what do we have ?

brandonHollinger:

14=-1x

mxddi3:

it's negative 14

brandonHollinger:

ok

mxddi3:

does that make sense so far?

brandonHollinger:

yeas

brandonHollinger:

yes

mxddi3:

ok so we have -14=-x now when you have a -x, how do you make it positive?

brandonHollinger:

i don't know.

mxddi3:

ok so when you have equations such as these \[-x=-14\] we have to divide BOTH sides by a -1 in order to make our x positive do you follow?

brandonHollinger:

yes

mxddi3:

ok so \[ \frac{ -x }{ -1 }= x\] then what is -14 divided by -1

brandonHollinger:

positive 14

mxddi3:

good, so we are left with x=14

mxddi3:

understand?

brandonHollinger:

yes thank you for your help

mxddi3:

you're welcome :)

brandonHollinger:

have a great evening.

mxddi3:

you too!